# GroundFlow: A Plug-in Module for Temporal Reasoning on 3D Point Cloud Sequential Grounding

### v1

@@ Abstract
Sequential grounding in 3D point clouds (SG3D) refers to locating sequences of objects by following text instructions for a daily activity with detailed steps. Current 3D visual grounding (3DVG) methods treat text instructions with multiple steps as a whole, without extracting useful temporal information from each step. However, the instructions in SG3D often contain pronouns such as “it”, “here” and “the same” to make language expressions concise. This requires grounding methods to understand the context and retrieve relevant information from previous steps to correctly locate object sequences. Due to the lack of an effective module for collecting related historical information, state-of-the-art 3DVG methods face significant challenges in adapting to the SG3D task. To fill this gap, we propose GroundFlow — a plug-in module for temporal reasoning on 3D point cloud sequential grounding. Firstly, we demonstrate that integrating GroundFlow improves the task accuracy of 3DVG baseline methods by a large margin (+7.5% and +10.2%) in the SG3D benchmark, even outperforming a 3D large language model pre-trained on various datasets. Furthermore, we selectively extract both short-term and long-term step information based on its relevance to the current instruction, enabling GroundFlow to take a comprehensive view of historical information and maintain its temporal understanding advantage as step counts increase. Overall, our work introduces temporal reasoning capabilities to existing 3DVG models and achieves state-of-the-art performance in the SG3D benchmark across five datasets.
@@ 3.1 Recurrent Framework with Temporal Fusion
As shown in Figure 2, our proposed recurrent framework takes the current step’s joint embedding $J_{t}$ and the previous step’s GroundFlow module output $\hat{J}_{t-1}$ as input, where the joint embedding with the hat ‘$\hat{\phantom{x}}$’ denotes the output after performing temporal fusion in GroundFlow. Consequently, the module’s output at the current step $\hat{J}_{t}$ serves as input for the next step’s inference. Since the joint embedding $J_{t}$ right before the grounding head is the only information required from 3DVG baselines to generate the corresponding $\hat{J}_{t}$ with important temporal information, our proposed recurrent framework is expected to adapt to a broader range of 3DVG model categories.
@@ 4.3 Comparison on SG3D Benchmark
Comparison with LLM-based methods. Following [55], an LLM with the 3D object classifier PointNet++ [36] is tested in a zero-shot manner. GPT-4 receives scene information, including each object’s ID, size, position, and the step instruction, while PointNet++, pre-trained on ScanNet, is used to predict the semantic labels for each object. The results in the first row of the table demonstrate poor zero-shot performance. Furthermore, the state-of-the-art 3D large language model, LEO, after fine-tuning on the SG3D benchmark is also compared. LEO, which uses Vicuna-7B [41] to process the token sequence, is pre-trained on an extensive range of 3D tasks, including object captioning [31, 54], object referring [51, 1, 20], 3D QA [17, 14, 50] and 3D navigation [29, 35], achieving top performance across various 3D point cloud benchmarks such as Scan2Cap [10], ScanQA [3] and SQA3D [32]. It is shown that LEO has decent results in SG3D benchmark and performs better than previous 3DVG methods due to its Internet-scale pre-trained knowledge.
@@ 4.4 Ablation Study
As shown in Table 2, GroundFlow demonstrates more effective improvements compared to classic temporal fusion methods. This advantage could stem from the limitations of existing methods: LSTM or GRU tends to forget long-term information. While transformers use attention mechanisms to capture long-range dependencies, they treat each step indiscriminately. As a result, they may still lose crucial historical information that is important at current step $t$ but not relevant at previous step $t-1$. Since previous step embeddings do not attend to this lost information, it cannot be carried forward to subsequent steps, even if it is essential for the current prediction. This issue is common in SG3D, where the current instruction often refers to objects from multiple past steps while intermediate instructions are irrelevant, as illustrated in the second example of Figure 5.
@@
Models Temporal Fusion Methods s-acc    t-acc $\Delta$s-acc $\Delta$t-acc
@@
Models Short-term Long-term s-acc   t-acc $\Delta$s-acc $\Delta$t-acc
@@ A SG3D Benchmark Statistics
Table 4: Dataset statistics of SG3D benchmark. #$P$, #$O$/$P$, #$T$ and #$S$ represent the number of 3D point cloud scenes, the average number of candidate objects per scene, the number of tasks and the number of steps, respectively.

### v3

@@ Abstract
Sequential grounding in 3D point clouds (SG3D) refers to locating sequences of objects by following text instructions for a daily activity with detailed steps. Current 3D visual grounding (3DVG) methods treat text instructions with multiple steps as a whole, without extracting useful temporal information from each step. However, the instructions in SG3D often contain pronouns such as “it”, “here” and “the same” to make language expressions concise. This requires grounding methods to understand the context and retrieve relevant information from previous steps to correctly locate object sequences. Due to the lack of an effective module for collecting related historical information, state-of-the-art 3DVG methods face significant challenges in adapting to the SG3D task. To fill this gap, we propose GroundFlow — a plug-in module for temporal reasoning on 3D point cloud sequential grounding. Firstly, we demonstrate that integrating GroundFlow improves the task accuracy of 3DVG baseline methods by a large margin (+7.5% and +10.2%) in the SG3D benchmark, even outperforming a 3D large language model pre-trained on various datasets. Furthermore, we selectively extract both short-term and long-term step information based on its relevance to the current instruction, enabling GroundFlow to take a comprehensive view of historical information and maintain its temporal understanding advantage as step counts increase. Overall, our work introduces temporal reasoning capabilities to existing 3DVG models and achieves state-of-the-art performance in the SG3D benchmark across five datasets. Codes are available at https://github.com/Jimntu/GroundFlow.
@@ 3.1 Recurrent Framework with Temporal Fusion
As shown in Figure 2, our proposed recurrent framework takes the current step’s joint embedding $J_{t}$ and the previous step’s GroundFlow module output $\hat{J}_{t-1}$ as input, where the joint embedding with the hat $\hat{\phantom{x}}$’ denotes the output after performing temporal fusion in GroundFlow. Consequently, the module’s output at the current step $\hat{J}_{t}$ serves as input for the next step’s inference. Since the joint embedding $J_{t}$ right before the grounding head is the only information required from 3DVG baselines to generate the corresponding $\hat{J}_{t}$ with important temporal information, our proposed recurrent framework is expected to adapt to a broader range of 3DVG model categories.
@@ 4.3 Comparison on SG3D Benchmark
Comparison with LLM-based methods. Following [55], an LLM with the 3D object classifier PointNet++ [36] is tested in a zero-shot manner. GPT-4 receives scene information, including each object’s ID, size, position, and the step instruction, while PointNet++, pre-trained on ScanNet, is used to predict the semantic labels for each object. The results in the first row of the table demonstrate poor zero-shot performance. Furthermore, the state-of-the-art 3D large language model, LEO, after fine-tuning on the SG3D benchmark is also compared. It is pre-trained on an extensive range of 3D tasks, including object captioning [31, 54], object referring [51, 1, 20], 3D QA [17, 14, 50] and 3D navigation [29, 35], achieving top performance across various 3D point cloud benchmarks. In fine-tuning stage, LEO predicts a special $GRD]_{t}$ token at each step $t$, which is concatenated with object tokens and passed to the grounding head to predict the target object $O_{t}$. It is shown that LEO has decent results in SG3D benchmark and performs better than previous 3DVG methods due to its Internet-scale pre-trained knowledge.
@@ 4.4 Ablation Study
As shown in Table 2, GroundFlow demonstrates more effective improvements compared to classic temporal fusion methods. This advantage could stem from the limitations of existing methods: LSTM or GRU tends to forget long-term information. While transformers use attention mechanisms to capture long-range dependencies, they treat each step indiscriminately. As a result, they may still lose crucial historical information that is important at current step $t$ but not relevant at previous step t1t-1. Since previous step embeddings do not attend to this lost information, it cannot be carried forward to subsequent steps, even if it is essential for the current prediction. This issue is common in SG3D, where the current instruction often refers to objects from multiple past steps while intermediate instructions are irrelevant, as illustrated in the second example of Figure 5.
@@
Models Temporal Fusion Methods s-acc    t-$\Deltas-acc Δ\Deltat$-acc
@@
Models Short-term Long-term s-acc   t-$\Deltas-acc Δ\Deltat$-acc
@@ 5 Conclusion
Acknowledgements
Zijun Lin is supported by the Agency for Science, Technology and Research (A*STAR) Computing and Information Science (ACIS) Scholarship. Shuting He is sponsored by Shanghai Pujiang Programme 24PJD030 and Natural Science Foundation of Shanghai 25ZR1402138. This research is supported in part by A*STAR SERC CRF funding to C.T., and in part by A*STAR IAF-ICP Programme I2501E0041 and the Schaeffler-NTU Corporate Lab (SHARE@NTU). The work was done at Rapid-Rich Object Search (ROSE) Lab, School of Electrical & Electronic Engineering, Nanyang Technological University.
@@ A SG3D Benchmark Statistics
Table 4: Dataset statistics of SG3D benchmark. #$P$, #OO/$P$, #$T$ and #$S$ represent the number of 3D point cloud scenes, the average number of candidate objects per scene, the number of tasks and the number of steps, respectively.